# L2R: Learning to Reduce Search Space for Generalizable Neural Routing Solver

### v1

@@ Appendix C Adaptation Attention Free Module
$\mathrm{Attention}=\sigma(Q)\odot\frac{\exp(A)(\exp(K)\odot V)}{\exp(A)\exp(K)},$ (23)
where $\sigma$ denotes the sigmoid function, $\odot$ represents the element-wise product, and $A=\{\mathbf{a}_{ij}\}$ denotes the pair-wise adaptation bias, it is calculated as:
$a_{ij}=-\alpha\cdot\log_{2}N\cdot d_{ij}\quad\forall i,j\in\{1,\ldots,N\},$ (24)
where $N$ is the total number of nodes, $d_{ij}$ represents the distance between node $i$ and node $j$, and $\alpha>0$ is a learnable parameter with a default value of 1. Compared to multi-head attention (MHA) (Vaswani et al., 2017), AAFM enables the model to explicitly capture instance-specific knowledge by updating pair-wise adaptation biases while exhibiting lower computational overhead. Further details are provided in the related work section mentioned above.
Appendix D Training
Following Kool et al. (2019), we employ an exponential baseline ($\beta=0.8$) during the first epoch to stabilize initial learning. The baseline parameters $\bm{\theta}^{BL}$ are updated only if the improvement is statistically significant, as determined by a paired t-test ($\alpha=5\%$) conducted on $10\text{,}000$ separate evaluation instances at the end of each epoch. If the baseline policy is updated, new evaluation instances are sampled to prevent overfitting. Additionally, at the first construction step for TSP, $\pi_{1}$ is randomly selected as the initial partial solution and is treated as both the first and last nodes for subsequent decoding steps.
@@ Appendix D Training
1:  Input: number of epochs $E$, steps per epoch $T$, batch size $B$, significance $\alpha$,
2:  Output: The trained model with parameters $\bm{\theta}=\{\bm{\theta}_{R},\bm{\theta}_{L}\}$, which are reduction model $\bm{\theta}_{R}$ and local construction model $\bm{\theta}_{L}$
3:  Init $\bm{\theta}_{R},\ \bm{\theta}_{R}^{BL}\leftarrow\bm{\theta}_{R}$ and $\bm{\theta}_{L},\ \bm{\theta}_{L}^{BL}\leftarrow\bm{\theta}_{L}$
@@
7:        $G_{i}^{{}^{\prime}}\leftarrow\text{StaticReduction()}\enspace\forall i\in\{1,\ldots,B\}$
8:        while not done do
9:           // $\mathcal{N}_{i,t}$ denotes the $k$ high-potential candidate nodes at step $t$ given sparse graph $G_{i}^{{}^{\prime}}$
10:           $\mathcal{N}_{i,t}\leftarrow$ DynamicReduction($G_{i}^{{}^{\prime}},\ \pi_{t-1}$$\bm{\theta}_{R}$) $\enspace\forall i\in\{1,\ldots,B\}$
11:           // using a decoding strategy of sampling according to the output probabilities
12:           $\pi_{i,t}\leftarrow$ LocalConstruction($\pi_{1},\ \pi_{t-1},\ \mathcal{N}_{i,t}$$\bm{\theta}_{L}$) $\enspace\forall i\in\{1,\ldots,B\}$
13:        end while
14:        while not done do
15:           // $\mathcal{N}_{i,t}^{BL}$ denotes the $k$ high-potential candidate nodes at step $t$ given sparse graph $G_{i}^{{}^{\prime}}$
16:           $\mathcal{N}_{i,t}^{BL}\leftarrow$ DynamicReduction($G_{i}^{{}^{\prime}},\ \pi_{t-1}^{BL}$$\bm{\theta}_{R}^{BL}$) $\enspace\forall i\in\{1,\ldots,B\}$
17:           // using a decoding strategy of greedy according to the output probabilities
18:           $\pi_{i,t}^{BL}\leftarrow$ LocalConstruction($\pi_{1}^{BL},\ \pi_{t-1}^{BL},\ \mathcal{N}_{i,t}^{BL}$$\bm{\theta}_{L}^{BL}$) $\enspace\forall i\in\{1,\ldots,B\}$
19:        end while
20:        $\nabla\mathcal{L}\leftarrow\sum_{i=1}^{B}\left(R(\bm{\pi}_{i})-R(\bm{\pi}_{i}^{BL})\right)\nabla_{\bm{\theta}}\log p_{\bm{\theta}}(\bm{\pi}_{i})$
21:        $\bm{\theta}\leftarrow\text{Adam}(\bm{\theta},\nabla\mathcal{L})$
22:     end for
@@
24:        $\bm{\theta}_{R}^{BL}\leftarrow\bm{\theta}_{R}$
25:        $\bm{\theta}_{L}^{BL}\leftarrow\bm{\theta}_{L}$
26:     end if
27:  end for
Appendix E Model Setting
Detailed information about the hyperparameter settings can be found in Table 7. Here, the number of layers denotes the number of attention layers of the local construction model.
Table 7: Model hyperparameter settings in experiments.
@@ Appendix E Model Setting
Initial learning rate $10^{-4}$
Learning rate decay $0.98$ per epoch
The number of attention layer $6$
Embedding dimension $128$
Feed forward dimension $512$
Clipping parameter 10
Training capacity -- $50$
Maximum search space size $k$ $20$ $50$
Percentage of static reduction $10\%$
Gradient clipping max_norm=$1.0$
Weight decay --
Batch size $180$ $64$
@@
Appendix F Detailed Results on Real-world Datasets
We evaluate the generalization performance on instances from CVRPLIB Set-XXL (Arnold et al., 2019) and TSPLIB (Reinelt, 1991). Detailed results for large-scale instances are provided in Table 8 and Table 9. The results demonstrate that L2R consistently outperforms other models across instances of varying scales, demonstrating its strong applicability in real-world scenarios.
Table 8: The detailed results on large-scale TSPLIB (Reinelt, 1991) instances ($9005,000\leq N\leq 85,$ 900).
Instance Scale BQ LEHD ELG INViT L2R
bs16 greedy aug×$\times$8 greedy greedy
rl5915 5,915 19.58% 24.17% OOM 14.02% 7.57%
rl5934 5,934 24.53% 24.11% OOM 12.91% 12.63%
pla7397 7,397 47.63% 40.94% OOM 9.45% 7.59%
rl11849 11,849 OOM 38.04% OOM 12.71% 7.93%
usa13509 13,509 OOM 71.10% OOM 13.44% 7.46%
brd14051 14,051 OOM 41.22% OOM 9.31% 6.53%
d15112 15,112 OOM 35.82% OOM 7.24% 6.60%
d18512 18,512 OOM OOM OOM 6.62% 5.23%
pla33810 33,810 OOM OOM OOM 7.04% 6.13%
pla85900 85,900 OOM OOM OOM 7.21% 4.62%
Sloved# 3/10 7/10 0/10 10/10 10/10
Avg.gap 30.58%† 39.34%† -- 10.00% 7.23%
@@ Appendix F Detailed Results on Real-world Datasets
Table 9: The detailed results on CVRPLIB XXL (Arnold et al., 2019) instances ($0003,000\leq N\leq 30,$ 000).
Instance Scale BQ LEHD ELG INViT L2R
bs16 greedy aug×$\times$8 greedy greedy
Leuven1 3,000 15.39% 16.60% 12.12% 13.71% 11.53%
Leuven2 4,000 25.69% 34.85% 21.52% 26.08% 16.14%
Antwerp1 6,000 13.64% 14.66% OOM 15.40% 13.14%
Antwerp2 7,000 26.09% 22.77% OOM 27.75% 12.63%
Ghent1 10,000 OOM 27.23% OOM 15.87% 10.29%
Ghent2 11,000 OOM 38.36% OOM 30.78% 12.78%
Brussels1 15,000 OOM OOM OOM 18.09% 12.93%
Brussels2 16,000 OOM OOM OOM 32.08% 12.61%
Flanders1 20,000 OOM OOM OOM 23.41% 7.81%
Flanders2 30,000 OOM OOM OOM 39.60% 13.05%
Sloved# 4/10 6/10 2/10 10/10 10/10
Avg.gap 20.20%† 25.75%† 16.82%† 24.28% 12.29%
  • Some instances are skipped due to the OOM issue.
Appendix G Ablation Study
G.1 Effects of Learning-based Reduction
@@
Table 10: Comparison between different reduction approaches on cross-distribution generalization with $5,000$ nodes. D-SSR denotes distance-based search space reduction.
Method Uniform Cluster Explosion Implosion
LKH3 0.00% 0.00% 0.00% 0.00%
D-SSR ($k=20$) 5.04% 5.73% 9.54% 6.20%
L2R ($k=20$) 4.62% 5.56% 8.78% 6.06%
G.2 L2R vs. Distance-based SSR with Larger Search Space
Table 11: Comparison between learning-based small space and distance-based large space on TSP instances with uniform distribution. D-SSR denotes distance-based search space reduction.
Method TSP1K TSP5K TSP10K TSP50K TSP100K
LKH3 0.00% 0.00% 0.00% 0.00% 0.00%
D-SSR ($k=50$) 5.51% 6.30% 6.73% 6.52% 6.62%
D-SSR ($k=75$) 5.13% 6.23% 6.95% 7.15% 7.33%
D-SSR ($k=100$) 4.50% 5.20% 5.39% 5.63% 5.74%
L2R ($k=20$) 4.42% 4.62% 4.71% 4.82% 4.70%
Appendix H Solution Visualizations
H.1 Solution Visualizations of Cross-distribution TSP Instances
(a) Optimal solution.
(b) L2R with greedy decoding (gap: 4.28%).
Figure 6: The solution visualizations of a TSP5K instance with uniform distribution.
(a) Optimal solution.
(b) L2R with greedy decoding (gap: 4.59%).
Figure 7: The solution visualizations of a TSP5K instance with cluster distribution.
(a) Optimal solution.
(b) L2R with greedy decoding (gap: 4.39%).
Figure 8: The solution visualizations of a TSP5K instance with explosion distribution.
(a) Optimal solution.
(b) L2R with greedy decoding (gap: 4.47%).
Figure 9: The solution visualizations of a TSP5K instance with implosion distribution.
H.2 Solution Visualizations of Large-scale TSPLIB Instances
(a) Instance nrw1379 (scale: 1379, gap: 3.39%).
(b) Instance u2319 (scale: 2319, gap: 0.43%).
(c) Instance pcb3038 (scale: 3038, gap: 4.89%).
(d) Instance fnl4461 (scale: 4461, gap: 4.28%).

### v2

@@ Appendix G Adaptation Attention Free Module
$\mathrm{Attention}(Q,K,V,A)=\sigma(Q)\odot\frac{\exp(A)(\exp(K)\odot V)}{\exp(A)\exp(K)},$ (36)
where $\sigma$ denotes the sigmoid function, $\odot$ represents the element-wise product, and $A=\{a_{ij}\}$ denotes the pair-wise adaptation bias. For more details about $a_{ij}$, please see Appendix D.
Compared to multi-head attention (MHA) [60], AAFM enables the model to explicitly capture instance-specific knowledge by updating pair-wise adaptation biases while exhibiting lower computational overhead. Further details are provided in the related work section mentioned above. For a detailed discussion of the comparison of AAFM and MHA in the L2R framework, please refer to Section L.3.
Appendix H Pseudocode for Training
Following Kool et al. [16], we employ an exponential baseline ($\beta=0.8$) during the first epoch to stabilize initial learning. The baseline parameters $\bm{\theta}^{BL}$ are updated only if the improvement is statistically significant, as determined by a paired t-test ($\alpha=5\%$) conducted on $10\text{,}000$ separate evaluation instances at the end of each epoch. If the baseline policy is updated, new evaluation instances are sampled to prevent overfitting. Additionally, at the first construction step for TSP, $\pi_{1}$ is randomly selected as the initial partial solution and is treated as both the initial and last nodes for the first decoding steps.
@@ Appendix H Pseudocode for Training
1:  Input: number of epochs $E$, steps per epoch $T$, batch size $B$, significance $\alpha$, the percentage $\gamma$ of static reduction.
2:  Output: The trained model with parameters $\bm{\theta}=\{\bm{\theta}_{R},\bm{\theta}_{L}\}$, which are reduction model $\bm{\theta}_{R}$ and local construction model $\bm{\theta}_{L}$.
3:  Init $\bm{\theta}_{R},\ \bm{\theta}_{R}^{BL}\leftarrow\bm{\theta}_{R}$ and $\bm{\theta}_{L},\ \bm{\theta}_{L}^{BL}\leftarrow\bm{\theta}_{L}$.
@@
7:        $G_{i}^{{}^{\prime}}\leftarrow\text{StaticReduction($\gamma$)}\enspace\forall i\in\{1,\ldots,B\}$
8:        // For $\bm{\theta}$, we use a decoding strategy of sampling according to the output probabilities.
9:        while not done do
10:           // $\mathcal{N}_{i,t}$ denotes the $k$ high-potential candidate nodes at step $t$ given sparse graph $G_{i}^{{}^{\prime}}$ generated by $\bm{\theta}_{R}$.
11:           // This selection $\tau_{i,t}$ is only used for the loss calculation of the reduction model.
12:           $\mathcal{N}_{i,t},\tau_{i,t}\leftarrow$ DynamicReduction($G_{i}^{{}^{\prime}},\ \pi_{t-1}$$\bm{\theta}_{R}$) $\enspace\forall i\in\{1,\ldots,B\}$
13:           // This selection $\pi_{i,t}$ generated by $\bm{\theta}_{L}$ is used as the node selection for the current step $t$.
14:           $\pi_{i,t}\leftarrow$ LocalConstruction($\pi_{1},\ \pi_{t-1},\ \mathcal{N}_{i,t}$$\bm{\theta}_{L}$) $\enspace\forall i\in\{1,\ldots,B\}$
15:        end while
16:        // For $\bm{\theta}^{BL}$, we use a greedy decoding strategy according to the output probabilities.
17:        while not done do
18:           $\mathcal{N}_{i,t}^{BL}\leftarrow$ DynamicReduction($G_{i}^{{}^{\prime}},\ \pi_{t-1}^{BL}$$\bm{\theta}_{R}^{BL}$) $\enspace\forall i\in\{1,\ldots,B\}$
19:           // For $\bm{\theta}^{BL}$, we only obtain the current node selection $\pi_{i,t}^{BL}$ of $\bm{\theta}_{L}^{BL}$ for calculating the reward.
20:           $\pi_{i,t}^{BL}\leftarrow$ LocalConstruction($\pi_{1}^{BL},\ \pi_{t-1}^{BL},\ \mathcal{N}_{i,t}^{BL}$$\bm{\theta}_{L}^{BL}$) $\enspace\forall i\in\{1,\ldots,B\}$
21:        end while
22:        $\nabla_{\bm{\theta}_{R}}\mathcal{L}_{R}(\bm{\theta}_{R})\leftarrow\frac{1}{B}\sum_{i=1}^{B}\left(\mathcal{R}(\pi_{i})-\mathcal{R}(\bm{\pi}_{i}^{BL})\right)\nabla_{\bm{\theta}_{R}}\log o(\tau_{i})$
23:        $\nabla_{\bm{\theta}_{L}}\mathcal{L}_{L}(\bm{\theta}_{L})\leftarrow\frac{1}{B}\sum_{i=1}^{B}\left(\mathcal{R}(\pi_{i})-\mathcal{R}(\bm{\pi}_{i}^{BL})\right)\nabla_{\bm{\theta}_{L}}\log p(\pi_{i})$
24:        $\mathcal{L}_{\mathrm{Joint}}(\bm{\theta})\leftarrow\mathcal{L}_{R}(\bm{\theta}_{R})+\mathcal{L}_{L}(\bm{\theta}_{L})$
25:        $\bm{\theta}\leftarrow\text{Adam}(\bm{\theta},\nabla\mathcal{L}_{\mathrm{Joint}})$
26:     end for
@@
28:        $\bm{\theta}_{R}^{BL}\leftarrow\bm{\theta}_{R}$
29:        $\bm{\theta}_{L}^{BL}\leftarrow\bm{\theta}_{L}$
30:     end if
31:  end for
Appendix I Training and Model Settings
Here, we provide the hyperparameter settings for the model and training process, and the detailed information can be found in Table 9.
Table 9: Training settings in our experiments.
@@ Appendix I Training and Model Settings
Initial learning rate of $\bm{\theta}_{R}$ $10^{-4}$
Learning rate decay of $\bm{\theta}_{R}$ $0.98$ per epoch
Initial learning rate of $\bm{\theta}_{L}$ $10^{-4}$
Learning rate decay of $\bm{\theta}_{L}$ $0.98$ per epoch
Batch size $180$ $60$
@@
The number of attention layer of $\bm{\theta}_{L}$ $6$
Embedding dimension $128$
Feed forward dimension $512$
Clipping parameter $\xi$ 10
Training capacity -- $50$
Maximum search space size $k$ $20$ $50$
Percentage $\gamma$ of static reduction $10\%$
Gradient clipping max_norm=$1.0$
Training baseline (epoch$\ =1$) Exponential baseline ($\beta=0.8$)
[…]
Appendix J Case-by-case Results on Benchmark Datasets
J.1 L2R vs. Algorithmic Methods
Taking the TSPLIB dataset as an example, we add the results of algorithmic methods (i.e., LKH3 and Concorde). As shown in Table 10, LKH3 and Concorde can outperform L2R on large-scale real-world instances, but they require a much longer runtime to obtain high-quality solutions. In addition, it should be emphasized that L2R can outperform all the other NCO methods with fast runtime on all comparisons, detailed results are provided in Table 11.
Table 10: Comparison between L2R and algorithmic methods (i.e., LKH-3 and Concorde) on large-scale TSPLIB instances ($9005,000\leq N\leq 85,$ 900). Here, "BKS" denotes the best-known solution.
Instance Scale BKS LKH-3 Concorde L2R Greedy L2R PRC100
Gap Time Gap Time Gap Time Gap Time
rl5915 5,915 565,530 0.00% 1.2h 0.29% 1.1h 10.04% 29s 6.37% 4.6m
rl5934 5,934 556,045 0.00% 1.3h 0.23% 1.2h 12.39% 29s 7.20% 5.2m
pla7397 7,397 23,260,728 0.00% 11.8h 0.22% 1.0h 7.22 % 36s 3.25% 5.2m
rl11849 11,849 923,288 0.00% 4.8h 0.66% 1.1h 6.96% 58s 4.49% 5.4m
usa13509 13,509 19,982,859 0.00% 7.5h 0.25% 1.2h 6.99% 1.1m 4.89% 5.4m
brd14051 14,051 469,385 0.00% 9.1h 0.39% 1.1h 6.52% 1.1m 4.16% 5.5m
d15112 15,112 1,573,084 0.00% 11.4h 0.26% 1.1h 7.55 % 1.2m 5.35% 5.2m
d18512 18,512 645,238 0.00% 13.1h 0.39% 1.1h 5.39% 1.5m 3.37% 5.8m
pla33810 33,810 66,048,945 0.00% 18h N/A 6.37% 2.8m 4.35% 7.4m
pla85900 85,900 142,382,641 0.00% 23h N/A 4.21% 7.7m 2.32% 12.9m
Avg.gap & Time 0.00% 10.1h 0.34%† 1.1h 7.36% 1.8m 4.58% 6.3m
  • The pla33810 and pla85900 instances are skipped due to infeasible solution generation.
J.2 L2R vs. Classical Neural Solvers
We evaluate the generalization performance on instances from CVRPLIB Set-XXL [35] and TSPLIB [29]. Detailed results for large-scale instances are provided in Table 11. The results demonstrate that L2R consistently outperforms other models across instances of varying scales, demonstrating its strong applicability in real-world scenarios.
Table 11: The detailed results on large-scale TSPLIB [29] instances ($9005,000\leq N\leq 85,$ 900) and CVRPLIB [35] instances ($0003,000\leq N\leq 30,$ 000).
Instance Scale GLOP BQ LEHD LEHD POMO ELG INViT L2R L2R
bs16 greedy RRC1000 aug×$\times$8 aug×$\times$8 greedy greedy PRC100
rl5915 5,915 11.56% 19.58% 24.17% 12.38% OOM OOM 14.02% 10.04% 6.37%
rl5934 5,934 11.32% 24.53% 24.11% 11.94% OOM OOM 12.91% 12.39% 7.20%
pla7397 7,397 6.38% 47.63% 40.94% 15.31% OOM OOM 9.45% 7.22% 3.25%
rl11849 11,849 9.55% OOM 38.04% 18.19% OOM OOM 12.71% 6.96% 4.49%
usa13509 13,509 5.81% OOM 71.10% 31.37% OOM OOM 13.44% 6.99% 4.89%
brd14051 14,051 4.87% OOM 41.22% 21.19% OOM OOM 9.31% 6.52% 4.16%
d15112 15,112 4.90% OOM 35.82% 18.83% OOM OOM 7.58% 7.55% 5.35%
d18512 18,512 4.99% OOM OOM OOM OOM OOM 6.62% 5.39% 3.37%
pla33810 33,810 9.53% OOM OOM OOM OOM OOM 7.04% 6.37% 4.35%
pla85900 85,900 7.97% OOM OOM OOM OOM OOM 7.21% 4.21% 2.32%
Sloved# 10/10 3/10 7/10 7/10 0/10 0/10 10/10 10/10 10/10
Avg.gap 7.69% 30.58%† 39.34%† 18.46%† -- -- 10.03% 7.36% 4.58%
Instance Scale GLOP- BQ LEHD LEHD POMO ELG INViT L2R L2R
LKH3 bs16 greedy RRC1000 aug×$\times$8 aug×$\times$8 greedy greedy PRC100
Leuven1 3,000 14.95% 15.39% 16.60% 10.71% 460.32% 12.12% 13.71% 11.49% 5.67%
Leuven2 4,000 16.54% 25.69% 34.85% 21.22% 202.17% 21.52% 26.08% 12.73% 11.82%
Antwerp1 6,000 19.08% 13.64% 14.66% 8.91% OOM OOM 15.40% 11.04% 5.93%
Antwerp2 7,000 17.72% 26.09% 22.77% 15.42% OOM OOM 27.75% 12.95% 11.07%
Ghent1 10,000 18.28% OOM 27.23% 17.28% OOM OOM 15.87% 10.15% 6.02%
Ghent2 11,000 16.45% OOM 38.36% 25.77% OOM OOM 30.78% 11.29% 11.01%
Brussels1 15,000 26.17% OOM OOM OOM OOM OOM 18.09% 12.65% 8.03%
Brussels2 16,000 17.56% OOM OOM OOM OOM OOM 32.08% 12.30% 11.79%
Flanders1 20,000 24.02% OOM OOM OOM OOM OOM 23.41% 7.85% 6.48%
Flanders2 30,000 25.42% OOM OOM OOM OOM OOM 39.60% 13.09% 12.70%
Sloved# 10/10 4/10 6/10 6/10 2/10 2/10 10/10 10/10 10/10
Avg.gap 19.62% 20.20%† 25.75%† 16.55%† 331.24%† 16.82%† 24.28% 11.55% 9.34%
@@
Appendix K Additional Analyses
K.1 L2R vs. D-SSR
To validate the effectiveness of our proposed L2R, we conduct a comparative experiment against a dynamic distance-based SSR model (i.e., D-SSR), ensuring identical experimental settings. We systematically analyze the advantages of L2R, mainly including: (1) overall performance on real-world instances, (2) optimality gap and optimality ratio per instance, (3) impacts of different candidate node selection strategies, and (4) solution visualizations comparing the two approaches.
K.1.1 Overall Performance on Real-world Instances
According to the results in Table 12, compared with traditional D-SSR, our proposed L2R can obtain better generalization performance in large-scale TSPLIB and CVRPLIB instances. Notably, this advantage becomes more pronounced when solving complex CVRP instances, where L2R significantly improves the final solution quality compared with D-SSR.
@@
Table 12: Optimality gap comparison between two different candidate node selection approaches. Here, D-SSR denotes distance-based search space reduction.
TSPLIB CVRPLIB-XXL
Method 1K $\leq N\leq$ 5K 5K <$N\leq$ 100K All 3K $\leq N\leq$ 7K 7K <$N\leq$ 30K All
(23 instances) (10 instances) (33 instances) (4 instances) (6 instances) (10 instances)
D-SSR greedy 9.65% 7.47% 8.98% 14.38% 12.19% 13.06%
L2R greedy 9.16% 7.36% 8.61% 12.05% 11.22% 11.55%
K.1.2 Optimality Gap vs. Optimality Ratio
We then measure the optimality gap and optimality ratio per instance. The optimality ratio is defined as the proportion of nodes where the optimal next-visit node is located within the $k$ candidate nodes of the current node. As demonstrated in Table 13, L2R achieves superior solution quality over D-SSR on real-world datasets. Although L2R exhibits only marginal improvement in the optimality ratio of next-visit node selection, the cumulative effect of over-pruning during the solution construction process leads to significant deviations from the optimal route in subsequent node selections, ultimately compromising the overall solution quality obtained by D-SSR.
Table 13: Comparison of different reduction approaches on TSPLIB and CVRPLIB instances. All results are obtained with greedy decoding. "#Accuracy" measures the number of construction steps where the optimal next-visit node is contained within the top-$k$ candidates selected by either learning-based or distance-based reduction methods. "Optimality Ratio" measures the proportion of nodes where the optimal next-visit node is located within the $k$ candidate nodes of the current node. (i.e., #Accuracy * 100% $/\ N$).
Instance Scale Gap $\downarrow$ #Accuracy (Optimality Ratio) $\uparrow$
D-SSR L2R $\Delta$Gap $\downarrow$ D-SSR L2R $\Delta$#Accuracy $\uparrow$
TSPLIB pcb1173 1173 5.17% 4.02% --1.15% 1062 (90.54%) 1067 (90.96%) +5
rl1304 1304 9.88% 8.14% --1.74% 1225 (93.94%) 1234 (94.63%) +9
rl1323 1323 16.44% 9.34% --7.10% 1242 (93.88%) 1252 (94.63%) +10
u1817 1817 17.60% 9.54% --8.06% 1614 (88.83%) 1630 (89.71%) +16
d2103 2103 20.50% 8.09% --12.41% 1988 (94.53%) 2000 (95.10%) +12
pr2392 2392 10.72% 8.08% --2.64% 2105 (88.00%) 2129 (89.01%) +24
fnl4461 4461 6.85% 4.22% --2.63% 3871 (86.77%) 3896 (87.34%) +25
rl11849 11849 8.13% 6.96% --1.17% 10851 (91.58%) 10864 (91.69%) +13
usa13509 13509 8.97% 6.99% --1.98% 11886 (87.99%) 11947 (88.44%) +61
pla33810 33810 7.57% 6.37% --1.20% 30263 (89.51%) 30463 (90.10%) +200
CVRPLIB Leuven1 3000 12.67% 11.49% --1.18% 2284 (76.13%) 2294 (76.47%) +10
Leuven2 4000 15.71% 12.73% --2.98% 3287 (82.18%) 3299 (82.48%) +12
Antwerp1 6000 13.12% 11.04% --2.08% 4573 (76.22%) 4597 (76.62%) +24
Antwerp2 7000 16.00% 12.95% --3.05% 5611 (80.16%) 5638 (80.54%) +27
Ghent1 10000 11.68% 10.15% --1.53% 7658 (76.58%) 7665 (76.65%) +7
Ghent2 11000 13.55% 11.29% --2.26% 9024 (82.04%) 9045 (82.23%) +21
Brussels2 16000 13.83% 12.30% --1.53% 13099 (81.87%) 13109 (81.93%) +10
K.1.3 Impacts of Candidate Node Selection Strategy
Figure 6 presents further analysis of candidate node selection across different reduction approaches at identical construction steps. The results in Figure 6(a)--Figure 6(b) indicate that while both D-SSR and L2R models successfully include the majority of optimal nodes within their generated neighborhoods, D-SSR exhibits significantly degraded solution quality. This degradation can be primarily attributed to its failure to include a critical minority of nodes within its candidate set. Consequently, the local construction model struggles in effectively selecting the optimal next-visit node. These critical nodes are exemplified by those to lead the sub-problem into a suboptimal state (see Figure 6(c)--Figure 6(d)) or manifest as anomalous single-node visits (see Figure 6(e)--Figure 6(f)). In contrast, the enhanced candidate selection capability of L2R allows it to capture more accurate candidate nodes, which facilitates better sequential decision-making and ultimately yields superior overall performance.
K.1.4 Solution Visualizations
To further support our argument, we provide additional solution visualizations comparing L2R and D-SSR. As illustrated in Figure 7, L2R tends to follow the optimal route more closely, but the cumulative effect of over-pruning during the solution construction of D-SSR leads to significant deviations in subsequent node selections. These deviations are a direct consequence of critical nodes not being adequately included in the candidate set, ultimately resulting in a significant degradation of D-SSR’s solution quality compared to L2R.
K.2 Complexity Analysis
We present the time and space complexity of three different constructive models: L2R with learning-based SSR, INViT with distance-based SSR, and LEHD without SSR. For the L2R method, we decompose the computational complexity into three key components: (1) the complexity of search space reduction (denoted as STEP1), (2) the complexity of local solution construction (denoted as STEP2), and (3) the overall framework complexity.
Table 14: Complexity analysis between L2R and existing works. Note that $N$ and $k$ denote the scale and the number of local neighbors, respectively. All results are obtained using greedy decoding. "Time" denotes the total inference time in the construction process, and "#$\#m$" is the number of tested instances.
Time Space TSP5K(#16) TSP10K(#16) CVRP5K(#100) CVRP10K(#16)
Method Complexity Complexity Gap Time Gap Time Gap Time Gap Time
LEHD $\mathcal{O}\left(N^{3}\right)$ $\mathcal{O}\left(N^{2}\right)$ 15.46% 24m 27.24% 3.1h 9.51% 2.3h 13.25% 3.2h
INViT-3V† $\mathcal{O}\left(N(k_{1}^{2}+k_{2}^{2}+k_{3}^{2}+k_{1})\right)$ $\mathcal{O}\left(k_{1}^{2}+k_{2}^{2}+k_{3}^{2}+k_{1}\right)$ 6.84% 3.4m 7.09% 9.3m 15.87% 11.6m 15.53% 19.2m
L2R-STEP1$\ddagger$ $\mathcal{O}\left(N^{2}\right)$ $\mathcal{O}\left(N\right)$- 8.0s -- 21.4s -- 17.5s -- 18.1s
L2R-STEP2$\ddagger$ $\mathcal{O}\left(Nk^{2}\right)$ $\mathcal{O}\left(k^{2}\right)$- 21.5s -- 42.6s -- 34.1s -- 52.1s
L2R $\mathcal{O}\left(N^{2}+Nk^{2}\right)$ $\mathcal{O}\left(N+k^{2}\right)$ 4.69% 29.5s 4.82% 1.1m 7.74% 51.6s 4.18% 1.2m
  • For INViT, $k_{1},k_{2},k_{3}$ represent different view, respectively.
  • $\ddagger$
    L2R-STEP1 and L2R-STEP2 represent the learning-based reduction model and local solution construction model, respectively.
Our comparative results in Table 14 demonstrate that L2R-STEP2 achieves significantly lower computational complexity than both LEHD and INViT. This efficiency advantage translates into substantially reduced runtime compared to existing representative models. Since distance-based SSR may compromise solution optimality, we introduce a learning-based reduction model to refine candidate node selection, thereby mitigating its negative impact. Although the reduction model incurs linear time complexity growth with problem size, experimental results demonstrate that the additional computational cost is justified, and L2R can perform better than D-SSR in solution quality. For more results and analyses, please see Section K.1.
(a) D-SSR (Gap:5.17%, #Optimal: 906)
(b) L2R (Gap:4.02%, #Optimal: 928)
(c) Partial solution and candidate nodes (D-SSR)
(d) Partial solution and candidate nodes (L2R)
(e) Partial solution and candidate nodes (D-SSR)
(f) Partial solution and candidate nodes (L2R)
Figure 6: The distributions of node selection strategy between different node selection approaches in TSPLIB, using pcb1173 as an example. The solutions are all generated by greedy decoding. In (a)-(b), the red connections and "#Optimal" indicate optimal node connections and the number of optimal connections, respectively. For the remaining figures, the red, green, and orange nodes indicate the current node, the next visiting node, and candidate nodes, respectively.
(a) Optimal solution (rl1323)
(b) D-SSR(Gap:16.44%,Ratio:93.88%)
(c) L2R(Gap:9.34%,Ratio:94.63%)
(d) Optimal solution (d2103)
(e) D-SSR(Gap:20.50%,Ratio:94.53%)
(f) L2R(Gap:8.09%,Ratio:95.10%)
(g) Optimal solution (fnl4461)
(h) D-SSR(Gap:6.85%,Ratio:86.77%)
(i) L2R(Gap:4.22%,Ratio:87.34%)